OpenStudy (anonymous):

can someone help explain how to determine if something is a vector space or not

OpenStudy (anonymous):

you have to check it for the axioms and see if they are true or not

OpenStudy (anonymous):

how do you use the axioms to determine for instance if the set of all fifth degree polynomials with the standard operations is a vestor space

OpenStudy (zarkon):

Is that space closed under addition

OpenStudy (anonymous):

i guess there is a subtlety here. if it means the polynomial must have degree 5, then there is no zero vector. if it means degree 5 or less, then yes

OpenStudy (zarkon):

right...as stated it is not a vector space...no zero vector and no closure of addition

OpenStudy (anonymous):

that is the very thing that Im confused about how do you determine that there is no zero vector and no closure of addition

OpenStudy (zarkon):

the zero vector has to be just the number 0 0 is not a 5th degree polynomial

OpenStudy (zarkon):

\[(x^5+3x)+(-x^5+x^2)=3x-x^2\] which is not a 5th degree polly

OpenStudy (anonymous):

is there a vector (fifth degree polynomial in this case) say \[p(x)\] with \[p(x)+v(x)=v(x)\] for all fifth degree polynomials v the answer is no, because the only polynomial that would work would be the zero polynomial, which does not have degree 5

OpenStudy (anonymous):

further more, \[x^5-2x+(-x^5)\] does not have degree 5, so it is not closed under addition

OpenStudy (anonymous):

what zarkon said

OpenStudy (anonymous):

ok that makes sense so what is the differance in saying it has fifth degree and fifth degree or less

OpenStudy (anonymous):

is one a vector and one is not and why

OpenStudy (zarkon):

5th degree or less is a vector space

OpenStudy (anonymous):

so what is the step by step method to determing the answer to these types of problems. I mean when talking about continuous functions and things like that wouldnt get a little tricky

OpenStudy (zarkon):

use the difinition I provided in my first post

OpenStudy (anonymous):

i know it has to adhere to the list of axioms but can you explain it a little.